For the reaction

3A(g)+2B(g)⇌C(g)

Kc = 73.0 at a temperature of 379 ∘C .

Calculate the value of Kp.

Express your answer numerically.

For the reaction

X(g)+2Y(g)⇌3Z(g)

Kp = 2.82×10−2 at a temperature of 195 ∘C .

Calculate the value of Kc.

Express your answer numerically.
Answer #1

3A(g)+2B(g)⇌C(g)

Kp = Kc (RT)dn

dn = = 1 - 5 = -4

Kc = 73.0

Kp = 73.0 (0.0821 x 652)-4

Kp = 73.0 (53.53)-4

Kp = 8.90 x 10-6

X(g)+2Y(g)⇌3Z(g)

Kp = Kc (RT)n

n = 3 - 3 = 0

so Kp =Kc

Kc = 2.82 x 10-2